OpenStudy (phebe):

MEDALSSS!!! Solve: a. √k – 8 = 28

OpenStudy (phebe):

@Troy.

OpenStudy (solomonzelman):

add 8 to both sides, and square both sides

OpenStudy (phebe):

ok so do i divide

OpenStudy (solomonzelman):

You add 8 to both sides first. Can you do that? What do you get ?

OpenStudy (phebe):

36

OpenStudy (solomonzelman):

yes !!

OpenStudy (phebe):

36k

OpenStudy (phebe):

so is that the answer @SolomonZelman

OpenStudy (solomonzelman):

\(\large\color{blue}{ \rm \sqrt{k}=36,~~~~~~~~k=?~}\)

OpenStudy (phebe):

do i put the square root of 36 @SolomonZelman

OpenStudy (solomonzelman):

You square both sides (raise both sides to the 2nd power)

OpenStudy (phebe):

so k to the 2 power and 36 to the secong power @SolomonZelman

OpenStudy (solomonzelman):

\(\large\color{blue}{ \rm \sqrt{k}=36~}\) \(\large\color{blue}{ \rm ( \sqrt{k})^2=(36)^2~}\) \(\large\color{purple }{ \rm k=1296}\)

OpenStudy (phebe):

u sure thas the answer

OpenStudy (solomonzelman):

Yes I am sure:)

OpenStudy (phebe):

thanx smart selff

OpenStudy (solomonzelman):

unless the question was \(\large\color{purple }{ \rm \sqrt{k-8}=1296}\)

OpenStudy (phebe):

no it w3asn't u were right at first

OpenStudy (solomonzelman):

Okay then 1296 is the ans.

OpenStudy (phebe):

ok thanx